# False Alarms, Real Damage: Adversarial Attacks Using LLM-based Models on Text-based Cyber Threat Intelligence Systems

### v1

@@ Abstract
Cyber Threat Intelligence (CTI) has emerged as a vital complementary approach that operates in the early phases of the cyber threat lifecycle. CTI involves collecting, processing, and analysing threat data to provide a more accurate and rapid understanding of cyber threats. Due to the large volume of data, automation through Machine Learning (ML) and Natural Language Processing (NLP) models is essential for effective CTI extraction. These automated systems leverage Open Source Intelligence (OSINT) from sources like social networks, forums, and blogs to identify Indicators of Compromise (IoCs). Although prior research has focused on adversarial attacks on specific ML models, this study expands the scope by investigating vulnerabilities within various components of the entire CTI pipeline and their susceptibility to adversarial attacks. These vulnerabilities arise because they ingest textual inputs from various open sources, including real and potentially fake content. We analyse three types of attacks against CTI pipelines - evasion, flooding, and poisoning- and assess their impact on the system’s information selection capabilities. Specifically, on fake text generation, the work demonstrates how adversarial text generation techniques can create fake cybersecurity and cybersecurity-like text that misleads classifiers, degrades performance, and disrupts system functionality. The focus is primarily on the evasion attack, as it precedes and enables flooding and poisoning attacks within the CTI pipeline. Our findings reveal that the False Positive Rate (FPR) for the evasion attack reached 97% for a specialized ML classifier model, indicating the model’s high vulnerability to adversarial samples. Additionally, an FPR of 75% is observed for ChatGPT-4o as a classifier, indicating its susceptibility to adversarial examples. These results underscore the need for an additional verification component at the early stage of the CTI pipeline to detect and filter out misinformation before it spreads through the system.
@@ 1 Introduction
The increasing complexity of cyber threats has made it difficult for existing security measures to provide effective protection [1]. Security systems such as firewalls and Intrusion Detection Systems (IDS) often cannot prevent breaches and fail to detect the spread of malicious activities. An enhancement is to keep defense mechanisms up to date with appropriate configurations while enriching them with relevant and up-to-date threat information. One promising approach to enhance defensive capabilities is Cyber Threat Intelligence (CTI), which provides insights into the tactics, techniques, and procedures (TTPs) used in modern cyberattacks, along with the corresponding detection and mitigation strategies. CTI is an emerging field that analyses trends in cybercrime, hacktivism, and cyber spying by utilizing diverse sources, including open-source intelligence (OSINT), social media, and human intelligence [2]. An important challenge for CTI systems is the potential for attackers to inject fake intelligence across OSINT sources such as Twitter (currently named X [^1] We use Twitter through the paper as our main dataset was published before the name change.), Stack Overflow, dark web forums, and blogs, which can compromise the reliability of the intelligence gathered and pose challenges for organizations using such sources.
@@
This work explores the behavior of CTI pipelines when data is collected from diverse and potentially unverified sources. This alternative approach avoids the limitations of relying solely on trusted sources and opens the possibility of exploring dark web forums. However, it becomes more vulnerable to misinformation and deliberate manipulation by adversaries. The CTI extraction tools’ failure to filter false or misleading information makes them vulnerable to adversarial exploitation. Attackers can manipulate these systems and compromise the integrity of the gathered threat intelligence.
@@
While prior work has emphasized the benefits of automation in CTI extraction by ML models, the impact of adversarial manipulation on data sources has received limited attention. This work focuses on this overlooked intersection, where automation interacts directly with deception. It assesses vulnerabilities of the CTI extraction pipeline rather than introducing entirely new attack methodologies, aiming to demonstrate how adversaries can create a platform to deploy fake data to systematically deceive CTI models and systems. Understanding these deception mechanisms allows for evaluating the robustness of automated threat intelligence extraction processes more effectively. In summary, the contributions of this work are as follows:
@@
    An integrated CTI extraction pipeline is proposed that captures and generalizes the typical stages and components used across prior CTI pipelines.
@@
    The analysis and investigation of the common limitations of CTI pipelines.
@@
    A comprehensive exploration of the threat landscape encountered by automated CTI systems, including attacks where adversaries intentionally induce FPs by injecting misleading information into the system.
@@
    An evaluation of the pipeline performance after applying evasion, flooding, and poisoning attacks.
@@ 2 Preliminary definitions
Real vs. Fake text. It is important to distinguish between real and fake input text in CTI systems, but there are no universally standardized definitions. Typically, text generated by a machine is labeled as fake, while human-written text is considered real [8]. However, this study adopts a task-specific definition: real text refers to genuine security-related information that reflects actual cyber threats or vulnerabilities; fake text refers to false or misleading information crafted to deceive either humans or automated systems. Such fakes may use cybersecurity-like terminology while not corresponding to real-world threats. For example, a real security-related tweet is: ”Vulnerability Details: CVE-2024-52046 (CVSS 10/10) Apache MINA Remote Code Execution (RCE) Vulnerability.” A fake version could be: ”Exploit released for CVE-2024-52046 allowing full control over Apache MINA servers. No patch available yet—act now to secure your systems!”. Although the fake version mimics the style and vocabulary of a real CTI statement, it conveys false urgency and misleading content.
Figure 1 illustrates the base taxonomy of text input used in this study. It categorizes text as either real or fake based on its truthfulness and intent. Furthermore, fake texts are subcategorized based on whether they are machine-generated or human-written.
@@ 3.2 AI-based analysis
Researchers have employed NLP techniques alongside ML models to extract CTI from textual data. Although the extraction methods vary depending on the type of information, the overall process of CTI extraction in the pipeline remains the same. This applies whether extracting IoCs, TTPs, or other relevant threat intelligence components. Figure 2 provides an overview of techniques commonly used in CTI extraction pipelines, including classification, clustering, Named Entity Recognition (NER), semantic role labeling, and relation extraction [9, 10, 11, 12, 6, 13, 14, 15, 16, 17]. These techniques enable models to identify patterns, extract key entities, and map relationships in text. Table I lists specific examples applied in various studies that use diverse data sources and describes the contribution of the techniques to the purpose of the CTI tasks. It is important to note that the implementation details and sequence of techniques may differ among studies based on the specific goals and datasets used.
@@ 3.3 Monitoring and validation
The monitoring stage comprises key components such as the dashboard, alarm validation, and structured format. The dashboard provides a visual overview of recent events, which helps security operations center (SOC) users maintain situational awareness and respond to threats, as supported by research [15] and tools such as [20]. These implementations underscore the importance of dashboards in centralizing information and supporting decision-making in cybersecurity operations. Alarm validation is a valuable task performed by a skilled human analyst who reviews and verifies alerts generated earlier in the pipeline [21]. The validated alerts are then organized and shared in structured formats such as STIX or TAXII [22, 23].
@@ 3.4 Threat scoring
This stage evaluates the severity and priority of detected threats [24]. Severity reflects the potential impact based on factors such as exploitability and system exposure, while priority defines the response order considering severity and likelihood of exploitation [25]. Priority helps allocate resources effectively and ensures a timely response [26]. Severity is typically labeled as Critical, High, Medium, or Low, whereas priority includes Urgent, ASAP, Within 24 hours, and Low Priority levels [27]. This distinction improves decision-making and response efficiency in SOCs.
@@ 3.5 Actionability
Feedback. Some CTI systems use feedback loops where SOC teams evaluate the quality of the extracted data, filtering out inaccuracies and improving future extractions by retraining models based on quality standards [31].
Attack group correlation. The extracted CTI helps link patterns between incidents with known threat actors and provides information on their tactics and priorities. This correlation helps to identify threats and allows defense contractors to prioritize responses based on the behavior and capabilities of the attacker [32].
@@ 4.1 Overview of attacks
Evasion attack. Evasion attacks occur during the test step, with trained models targeted to be misled. Once the extractor pipeline is trained, adversaries can generate Fake Negative (FaN) input texts to fool the model into misclassifying them as positives. This manipulation leads to increased FP and FN texts appearing on the dashboard, confusing SOC professionals and diverting attention from real threats [33]. In the test scenario, this attack focuses on generating FaN inputs that resemble real cybersecurity texts but are intentionally designed to cause misclassifications. For example, for a binary classifier that distinguishes between security- and non-security-related texts [14], an adversary can lead the model to misclassify non-security texts as security-related, possibly disrupting decision-making.
This attack is expressed as $\text{argmax}_{\hat{x}}\left\{\text{Loss}\left(f\left(\hat{x}\right),y\right)\right\}$, where $f\left(\hat{x}\right)$ is the classifier, $\hat{x}$ is the modified input, and $y$ represents the true label. The goal is to modify inputs $x$ to maximize the loss function, leading the model to incorrect predictions. This indicates how small targeted changes to the input can result in misclassification, which is central to evasion attacks. Evasion attacks can be categorized into two main approaches: maximum-confidence and minimum-distance [34]. Maximum-confidence attacks aim to create adversarial examples that are misclassified with high confidence, but often involve substantial changes to the input that decrease its similarity. In contrast, minimum-distance attacks focus on minor and subtle modifications to maintain high similarity and make detection more difficult.
@@
Flooding attack. Flooding attacks [35] overwhelm the CTI extraction pipeline by injecting a high volume of deceptive FaN and Fake Positive (FaP) texts into the system dashboard, making it difficult for security analysts to distinguish between real and misleading alerts. Additionally, attackers may use this technique to conceal their malicious activities within the flood of data to evade detection. This strategy can be considered a form of denial-of-service (DoS) attack, where the system becomes overwhelmed with excessive traffic, preventing it from processing legitimate requests. As a result, real positives become obscured, diminishing the system’s ability to handle real threats. This overload weakens the reliability of the threat intelligence pipeline and increases ambiguity in threat detection, making it harder for analysts to respond to genuine risks. Beyond disrupting decision-making, flooding attacks waste the analyst’s effort, leading to system failure and degrading the overall security response efficiency. Repeatedly encountering false alerts results in a loss of trust in the model’s predictions, reducing confidence in automated classification outcomes. Furthermore, overwhelming SOC analysts strains security operations and effectively impairs their ability to manage real threats. The excessive volume of injected data also increases costs owing to log storage, requiring more computational resources to process and archive irrelevant information. This not only affects infrastructure overhead but also directly impacts the analyser’s ability to process and interpret incoming data accurately. When encountering fake inputs, the analyser has two possible reactions: discard them, wasting computational resources and analyst time; or misclassify them as genuine, leading to a poisoning attack.
@@ 4.2 Attacker conceptual model
Attacker’s capability. In the evasion attack, the attacker generates adversarial inputs with minimal detectable changes to exploit model vulnerabilities and degrade performance. In the flooding attack, the attacker injects coherent false data streams to overwhelm the system, disrupting data integrity and operational efficiency. Additionally, in a poisoning attack, the attacker generates FaN texts that closely resemble real security-related texts, deceive both the ML models and analysts, and ultimately corrupt the training dataset.
@@ 5.1 Dataset
A publicly available Twitter dataset with 31281 tweets [39] was the basis for generating FaN and FaP texts. Although it was collected in 2016, it offers unique advantages that make it suitable for this study. Since the target models are binary classifiers, a dataset with a binary classification structure is essential. It was employed to train the target model [14], ensuring experimental consistency and compatibility with the attack scenario. It includes manually labeled named entities extracted from the cybersecurity-related tweets, essential for generating realistic FaP texts in the flooding attack setup. The dataset includes a classification label (”relevant class”) that indicates whether a tweet is security-related (1) or not (0). Although the tweet sizes of the dataset are limited to 256 characters, the FaP and FaN texts generated for evasion, flooding, and poisoning attacks vary in length. The clean_tweet feature in the dataset, which contains refined tweet text, was used to create adversarial examples.

### v2

@@ Abstract
Cyber Threat Intelligence (CTI) has emerged as a vital complementary approach that operates in the early phases of the cyber threat lifecycle. CTI involves collecting, processing, and analysing threat data to provide a more accurate and rapid understanding of cyber threats. Due to the large volume of data, automation through Machine Learning (ML) and Natural Language Processing (NLP) models is essential for effective CTI extraction. These automated systems leverage Open Source Intelligence (OSINT) from sources like social networks, forums, and blogs to identify Indicators of Compromise (IoCs). Although prior research has focused on adversarial attacks on specific ML models, this study expands the scope by investigating vulnerabilities within various components of the entire CTI pipeline and their susceptibility to adversarial attacks. It particularly focuses on a system-level analysis of how existing adversarial techniques interact with and across multiple stages of this pipeline, resulting in cascading attack effects. These vulnerabilities arise because they ingest textual inputs from various open sources, including real and potentially fake content. We analyse three types of attacks against CTI pipelines - evasion, flooding, and poisoning and assess their impact on the system’s information selection capabilities Specifically, focusing on fake text generation, the work demonstrates how adversarial text generation techniques can create fake cybersecurity and cybersecurity-like text that misleads classifiers, degrades performance, and disrupts system functionality. The focus is primarily on the evasion attack, as it precedes and enables flooding and poisoning attacks within the CTI pipeline. Our findings reveal that the False Positive Rate (FPR) for the evasion attack reached 97% for a specialized ML classifier model, indicating the model’s high vulnerability to adversarial samples. Additionally, an FPR of 75% is observed for ChatGPT-4o as a classifier, indicating its susceptibility to adversarial examples. These results underscore the need for an additional verification component at the early stage of the CTI pipeline to detect and filter out misinformation before it spreads through the system.
@@ 1 Introduction
The increasing complexity of cyber threats has made it difficult for existing security measures to provide effective protection [47]. Security systems such as firewalls and Intrusion Detection Systems (IDS) often cannot prevent breaches and fail to detect the spread of malicious activities. An enhancement is to keep defense mechanisms up to date with appropriate configurations while enriching them with relevant and up-to-date threat information. One promising approach to enhance defensive capabilities is Cyber Threat Intelligence (CTI), which provides insights into the tactics, techniques, and procedures (TTPs) used in modern cyberattacks, along with the corresponding detection and mitigation strategies. CTI is an emerging field that analyses trends in cybercrime, hacktivism, and cyber spying by utilizing diverse sources, including open-source intelligence (OSINT), social media, and human intelligence [55]. An important challenge for CTI systems is the potential for attackers to inject fake intelligence across OSINT sources such as Twitter (currently named X [^1] We use Twitter through the paper as our main dataset was published before the name change.), Stack Overflow, dark web forums, and blogs, which can compromise the reliability of the intelligence gathered and pose challenges for organizations using such sources. Although adversary-driven poisoning of CTI platforms (as in a confirmed incident) is rarely disclosed publicly, there are highly relevant cases of misinformation, fabricated cyber narratives, and deception that affected CTI platforms [64]. Examples of the dissemination of fake CTI reports regarding ransomware, governmental organizations, and large companies have been reported [37]. For instance, threat actors and scammers on web forums have repeatedly claimed major data breaches or offered “fresh” stolen datasets that turn out to be fake or repackaged old leaks. These false breach narratives often get re-reported by media or aggregated into CTI feeds before verification, causing wasted defensive effort and misprioritization [1].
@@
This work explores the behavior of CTI pipelines when data is collected from diverse and potentially unverified sources. This alternative approach avoids the limitations of relying solely on trusted sources and opens the possibility of exploring dark web forums. However, it becomes more vulnerable to misinformation and deliberate manipulation by adversaries. The CTI extraction tools’ failure to filter false or misleading information makes them vulnerable to adversarial exploitation. Attackers can manipulate these systems and compromise the integrity of the threat intelligence gathered.
@@
While prior work has emphasized the benefits of automation in CTI extraction by ML models, the impact of adversarial manipulation on data sources has received limited attention. Most existing CTI security research focuses on model-centric defences [Sarker2023Multi‐aspects], emphasizing classifier robustness, entity extraction accuracy, or isolated defense mechanisms, and implicitly assumes that data sources are trustworthy [53, 4, 41]. However, the systemic risks introduced by adversarial text generation at the data ingestion level and its compound effect on CTI pipelines remains insufficiently studied. This work focuses on this overlooked intersection, where automation interacts directly with deception. It assesses the vulnerabilities of the CTI extraction pipeline, rather than introducing entirely new attack methodologies, aiming to demonstrate how adversaries can create a platform to deploy fake data and systematically deceive CTI models and systems. Besides the fake text generation methodology, the novelty of this study lies in a system-level examination of the cascading effects caused by existing adversarial text generation techniques across the CTI pipeline. Understanding these deception mechanisms allows for evaluating the robustness of automated threat intelligence extraction processes more effectively. In summary, the contributions of this work are as follows:
@@
    An integrated CTI extraction pipeline model is proposed that captures and generalizes the typical stages and components used across prior CTI pipelines.
@@
    An attention-based fake text generation methodology using LLM chatbots and prompt optimisation.
@@
    System-level analysis of structural vulnerabilities in CTI pipelines, focusing on how weaknesses arise from interactions between the interconnected stages.
@@
    A comprehensive characterization of adversarial threats targeting automated CTI systems, including scenarios where attackers intentionally induce false positives by injecting misleading intelligence into data sources.
  5. 5.
    An evaluation of successive adversarial effects by applying evasion, flooding, and poisoning attacks and analyzing their combined impact on the pipeline performance.
Each contribution addresses a distinct limitation in the existing CTI security research. Together, they provide a unified analysis of adversarial risks on a CTI pipeline model that are often examined separately. By shifting the focus from individual models to interactions across the whole CTI pipeline, this study demonstrates how adversarial effects can accumulate over successive stages. The fake text generation methodology describes realistic misinformation approaches that rely on generative chatbot LLM models rather than constrained perturbations. Finally, the evaluation demonstrates that evasion attacks enable downstream flooding and poisoning, revealing system-level vulnerabilities that are not visible in prior work model-centric analysis.
@@ 2 Preliminary definitions
Real vs. Fake text. It is important to distinguish between real and fake input text in CTI systems, but there are no universally standardized definitions. Typically, text generated by a machine is labeled as fake, while human-written text is considered real [17]. However, this study adopts a task-specific definition: real text refers to genuine security-related information that reflects actual cyber threats or vulnerabilities; fake text refers to false, misleading, or manipulated cybersecurity information crafted to deceive either humans or automated systems. Such fakes may use cybersecurity-like terminology while not corresponding to real-world threats. For example, a real security-related tweet is: ”Vulnerability Details: CVE-2024-52046 (CVSS 10/10) Apache MINA Remote Code Execution (RCE) Vulnerability.” A fake version could be: ”Exploit released for CVE-2024-52046 allowing full control over Apache MINA servers. No patch available yet—act now to secure your systems!”. Although the fake version mimics the style and vocabulary of a real CTI statement, it conveys a false sense of urgency and misleading content.
Figure 1 illustrates the taxonomy of the text input used in this study. First, the text is classified as related or unrelated to cybersecurity. Then, the taxonomy distinguishes fake texts based on whether it is a human-written or machine-generated text. Finally, fake machine-generated texts divides into two categorises including AI-generated or Rule-based.
@@ 3.2 AI-based analysis
Researchers have employed NLP techniques alongside ML models to extract CTI from textual data. Although extraction methods vary depending on the type of information, the overall CTI extraction process in the pipeline remains the same. This applies whether extracting IoCs, TTPs, or other relevant threat intelligence components. Figure 2 provides an overview of techniques commonly used in CTI extraction pipelines, including classification, clustering, Named Entity Recognition (NER), semantic role labeling, and relation extraction [62, 5, 29, 57, 69, 50, 21, 9, 42, 39]. These techniques enable models to identify patterns, extract key entities, and map relationships in text. Table I lists specific examples applied in various studies that use diverse data sources and describes the contribution of the techniques to the purpose of the CTI tasks. It is important to note that the implementation details and the sequence of techniques may differ across studies, depending on the specific goals and datasets used.
@@ 3.3 Monitoring and validation
The monitoring stage comprises key components such as the dashboard, alarm validation, and a structured format. The dashboard provides a visual overview of recent events, which helps security operations center (SOC) users maintain situational awareness and respond to threats, as supported by research [9] and tools such as [23]. These implementations underscore the importance of dashboards in centralizing information and supporting decision-making in cybersecurity operations. Alarm validation is a valuable task performed by a skilled human analyst who reviews and verifies alerts generated earlier in the pipeline [2]. The validated alerts are then organized and shared in structured formats such as STIX or TAXII [48, 13].
@@ 3.4 Threat scoring
This stage evaluates the severity and priority of detected threats [6]. Severity reflects the potential impact based on factors such as exploitability and system exposure, while priority defines the response order based on severity and the likelihood of exploitation [31]. Priority helps allocate resources effectively and ensures a timely response [28]. Severity is typically labeled as Critical, High, Medium, or Low, whereas priority includes Urgent, ASAP, Within 24 hours, and Low Priority levels [66]. This distinction improves decision-making and response efficiency in SOCs.
@@ 3.5 Actionability
Feedback. Some CTI systems use feedback loops in which SOC teams evaluate the quality of extracted data, filtering out inaccuracies and improving future extractions by retraining models against quality standards [71].
Attack group correlation. The extracted CTI helps link patterns between incidents with known threat actors and provides information on their tactics and priorities. This correlation helps identify threats and allows defense contractors to prioritize responses based on the attacker’s behavior and capabilities [36].
@@ 4.1 Overview of attacks
Evasion attack. Evasion attacks occur during the test step, with trained models targeted to be misled. Once the extractor pipeline is trained, adversaries can generate Fake true Negative (FaN) input texts to fool the model into misclassifying them as positives. This manipulation leads to increased FP and FN texts appearing on the dashboard, confusing SOC professionals and diverting attention from real threats [46]. In the test scenario, this attack focuses on generating FaN inputs that resemble real cybersecurity texts but are intentionally designed to cause misclassifications. For example, for a binary classifier that distinguishes between security- and non-security-related texts [21], an adversary can lead the model to misclassify non-security texts as security-related, possibly disrupting decision-making. This attack is expressed as,
[…]
where f(x^)f($\hat${x}) indicates a binary classifier that predicts whether an input text is security-related or non-security-related, $\hat{x}$ represents the adversarially modified input text, and $y\in\{0,1\}$ refers to the ground-truth class label (0: non-security-related, 1: security-related). In this work, the loss function is the binary cross-entropy. The adversary seeks to modify the original input $x$ into $\hat{x}$ to maximize the classification loss, thereby increasing the likelihood of misclassification at inference time. Evasion attacks can be categorized into two main approaches: maximum-confidence and minimum-distance [35]. Maximum-confidence attacks aim to create adversarial examples that are misclassified with high confidence, but often involve substantial changes to the input that decrease its similarity. In contrast, minimum-distance attacks focus on minor and subtle modifications to maintain high similarity and make detection more difficult.
@@
Flooding attack. To implement our flooding attack, we need to generate manipulated real cybersecurity texts known as Fake true Positives (FaP). Flooding attacks [56] overwhelm the CTI extraction pipeline by injecting a high volume of deceptive FaN and FaP texts into the system dashboard, making it difficult for security analysts to distinguish between real and misleading alerts. Additionally, attackers may use this technique to conceal their malicious activities within the flood of data to evade detection. This strategy can be considered a form of denial-of-service (DoS) attack, where the system becomes overwhelmed with excessive traffic, preventing it from processing legitimate requests. As a result, real positives become obscured, diminishing the system’s ability to handle real threats. This overload weakens the reliability of the threat intelligence pipeline and increases ambiguity in threat detection, making it harder for analysts to respond to genuine risks. Beyond disrupting decision-making, flooding attacks waste the analyst’s effort, leading to system failure and degrading the overall security response efficiency. Repeatedly encountering false alerts results in a loss of trust in the model’s predictions, reducing confidence in automated classification outcomes. Furthermore, overwhelming SOC analysts strains security operations and effectively impairs their ability to manage real threats. The excessive volume of injected data also increases costs owing to log storage, requiring more computational resources to process and archive irrelevant information. This not only affects infrastructure overhead but also directly impacts the analyser’s ability to process and interpret incoming data accurately. When encountering fake inputs, the analyser has two possible reactions: discard them, wasting computational resources and analyst time; or misclassify them as genuine, leading to a poisoning attack.
@@ 4.2 Attacker conceptual model
Attacker’s capability. We explain the attacker capability requirements, classifying attacks by increasing thresholds of required resources, persistence, and influence over the CTI pipeline. Specifically, we distinguish three attacker capability levels: low-capability, mid-capability, and high-capability.
In the evasion attack, the attacker generates adversarial inputs with minimal detectable changes to exploit model vulnerabilities at inference time. This low-capability attack can be conducted at inference time using publicly accessible tools, does not require access to model internals, and requires the least attacker capability.
In the flooding attack, the attacker injects coherent false data streams to overwhelm the system, disrupting data integrity and operational efficiency. Such mid-capability attacks require moderate computational resources and sustained activity to continuously flood the data ingestion process. Additionally, in a poisoning attack, the attacker generates FaN texts that closely resemble real security-related texts, deceive both the ML models and analysts, and ultimately corrupt the training dataset. This is a high-capability attack because it requires persistence over time, computationally optimised text generation, and the ability to generate diversified FaN texts that will pass the human analysts, reach the training data, and influence the model decision boundary.
@@ 5.1 Dataset
A publicly available Twitter dataset with 31281 tweets [20] was the basis for generating FaN and FaP texts ,of which 11074 (35.4%) are security-related and the remaining 20207 (64.6%) are non-security tweets. Although the dataset was originally collected across three time periods (2016, 2017, and 2018), it was later annotated for NER and republished by Dionisio et al. [20]. This dataset offers unique advantages that make it suitable for this study. It was employed to train the target model [21], ensuring experimental consistency and compatibility with the attack scenario. It includes human-annotated named entities extracted from the cybersecurity-related tweets, essential for generating realistic FaP texts in the flooding attack setup. To the best of our knowledge, no other publicly available CTI corpus simultaneously provides binary relevance labels and manually annotated named entities, both of which are required for designing our flooding and poisoning attacks.
The dataset includes a classification label (”relevant class”) that indicates whether a tweet is security-related (1) or not (0). Although the tweet sizes of the dataset are limited to 256 characters, the FaP and FaN texts generated for evasion, flooding, and poisoning attacks vary in length. The clean_tweet feature in the dataset, which contains refined tweet text, was used to create adversarial examples. In addition to the Twitter dataset, we evaluated our approach using the AnnoCTR [33] dataset of 400 CTI reports, which consists entirely of cybersecurity-related content collected between 2020 and 2024. The longer text lengths in AnnoCTR enabled the generation of more elaborate and detailed FaN texts. From the dataset, which contains cyber threat reports segmented into individual text fragments, we sampled 2000 segments. Of these, for comparison purposes, 1500 were selected randomly to approximate the length distribution of Twitter posts (100-256 characters), while the remaining 500 consist of longer segments ranging from 256 to 1000 characters.